UW CHEMISTRY NMR

Announcments & Equipment Logs

March, 2016

UPDATE ! DRX499 – WHAT WENT WRONG WITH THE PROBE ?

First of, the good news is that 13C channel of the probe is made to work normally after fixing the following problem.  Here is a snap shot of the wobb : The 13C channel tune was broken, most probably because of excessive force used on the 13C tuning rod.   We have been appealing to the […]

Posted in Hardware, Spectrometer, Usage | No Comments »

ALERT ! DRX499 – ISSUES ON 13C CHANNEL

We are aware of the tuning problems on 13C channel of the TXO probe and it requires a careful look inside the probehead.  We will do this coming Thursday the 17th forenoon.  If the fix is not a quick one, we will be placing the BBO-Z gradient probe in its place, at least temporarily.  Please […]

Posted in Hardware, Spectrometer | No Comments »

UPDATE – AV500 – VT SYSTEM OK !

It turned out that there were lose connections in the cables to the VT sensor as well as heater in the probehead, that led to observed symptoms. These were sorted out and the VT system is functioning ok now. please proceed to setup your temperatures as usual.

Posted in Hardware, Spectrometer, Usage | No Comments »

ALERT ! AV500 – APPARENT VT SYSTEM ISSUES

Several users have posted in the log book about the the VT sensor in the probe not working as expected, leading to absence of temperature control.  Users are requested to run their experiment at room temperature for the present. We will be do a thorough investigation tomorrow beginning 10 am and post relevant updates on […]

Posted in Hardware, Spectrometer | No Comments »

UPDATE ! AV500 — SYSTEM IS FULLY OPERATIONAL

Users can start booking time on the instrument immediately. The probehead is the same as before i.e. BBI – Z-gradient.  The issue with this probe has been resolved. Topshim has been tested and working correctly both in 1D and 3D modes. Please contact Adrienne or myself, if you have any questions.   Sorry of this long […]

Posted in Hardware, Spectrometer, Usage | No Comments »

ALERT ! AV500 – SYSTEM DOWN !…

Being in Seattle, I am sure you can identify, when I say “When it rains it pours ! “ Right after the replacement of the faulty circuit boards, we had a short power disruption yesterday and it took some reconfiguration in the console to bring things to normal. Today, I have found the BBI probehead, […]

Posted in Hardware, Spectrometer | No Comments »

ALERT ! AV500 – SYSTEM FUNCTIONAL WITH A POSSIBLE CAVEAT

The power fluctuation yesterday left a few issues lurking in the console, which are resolved now.   The X-channel is working as expected but the 1H channel still has a detuning problem.   Due to this you will notice rather noisy 1H spectrum – this is related to the wobb curve simply moving to a detuned state […]

Posted in Uncategorized | No Comments »

UPDATE – AV301 : VT SYSTEM IN STABLE MODE

As mentioned, we have isolated the chiller that was freezing up samples when the probe heater was turning off.  You can collect data at room temperature and above currently.  We will update you when the low T capability is restored.

Posted in Hardware, Spectrometer | No Comments »

ALERT AV500 – 1H WOBB INOPERABLE

After the short power outage at approximately 3 pm on March 1, the av500 is not communicating with one of the new boards. NMR Facility managers will take a look at it first thing Wednesday morning. In the meantime, please reschedule your av500 time to another instrument.

Posted in Hardware, Spectrometer | No Comments »

Item Categories